How many mL of 20% acid solution must a scientist add to 80 mL of 5% solution to create a 10% acid solution?
Answer by CharlesG2(834)

How many mL of 20% acid solution must a scientist add to 80 mL of 5% solution to create a 10% acid solution?
amount A * %A + amount B * %B = amount C * %C
x * 0.20 + 80 * 0.05 = (x + 80) * 0.10
0.20x + 4 = 0.10x + 8
0.10x = 4
x = 40 ml